OpenStudy (anonymous):

verify the identity 1 + (sec^2)x *(sin^2)x = (sec^2)x

OpenStudy (shamim):

write something @kbarclay

OpenStudy (anonymous):

can you substitute the one as cos^2x/cos^2x ?

OpenStudy (shamim):

ya

OpenStudy (unklerhaukus):

times both sides by cos^2. and simplify

OpenStudy (shamim):

u know\[\sec x=\frac{ 1 }{ \cos x }\]

OpenStudy (anonymous):

yea so if I do that then cos^2x/ cos^2x + sin^2x/ cos^2x = sec^2x which then can it be cos^2x +sin^2x/ cos^2x ?

OpenStudy (raden):

sec^2 x * sin^2 x = sin^2 x/cos^2 x = tan^2 x see your left side becomes, 1 + tan^2 x, this is an identity in trigo, that's is sec^2 x LHS = RHS

OpenStudy (anonymous):

thanks

OpenStudy (unklerhaukus):

\[1 + \sec^2x \times\sin^2x = \sec^2x\] \[\qquad\qquad\qquad1 + \frac1{\cos^2x} \times\sin^2x = \frac1{\cos^2x}\\ ~~~\cos^2x\times\left(1 + \frac1{\cos^2x} \times\sin^2x\right) = \cos^2x\times\frac1{\cos^2x}\\\qquad\qquad\qquad\qquad\qquad\quad\qquad=\]

OpenStudy (anonymous):

wait how does that simplify to sec^2x=sec^2x @UnkleRhaukus

OpenStudy (unklerhaukus):

if you do it my way it should simplify to cos^2+sin^2=1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so that would make it 1=1 since cos^2x+sin^2x = 1 and cos^2x/cos^2x =1

OpenStudy (unklerhaukus):

well yeah,

OpenStudy (unklerhaukus):

\[\boxed{\color{seagreen}{\mathcal{Welcome~to~\color{skyblue}{OpenStudy}}}\\\qquad\color{orangered}{\ast\text{kbarclay}\ast}}\]